OpenStudy (anonymous):

If each of two complementary angles has the same measure, then each angle will equal _____. 180° 90° 45° 22.5°

OpenStudy (anonymous):

90 degrees cause 90 + 90 is 180, which is comp.

OpenStudy (anonymous):

so its 90

OpenStudy (unklerhaukus):

complementary angles add to a right angle A+B=90° if both angles are the same A+A=90°

jimthompson5910 (jim_thompson5910):

if x and y are complementary, then x+y = 90 if they have the same measure, then y = x and you can say x+x = 90 so x = ???

OpenStudy (anonymous):

ok thanks so it 45?

OpenStudy (unklerhaukus):

\[\checkmark\]

OpenStudy (anonymous):

thanks!!!